# Neural Scaling Laws on Graphs

### v1

@@ Appendix F The Relationship between Edge Number and FLOPs
where $U$ and $M$ can be assumed to contain $x$ and $y$ add-multiply operations, respectively. Suppose the total node number in the dataset is N, the total edge number is E, and the model has L layers. For node $i$, aggregating messages from its neighbors needs $x+yd_{i}$ add-multiply operations, where $d_{i}$ is the degree of node $i$. For the whole datasets, there would be $\sum_{i}x+yd_{i}\approx 2yE$. For the whole model, one feed-forward pass needs $2yEL$ add-multiply operations. Thus, we have shown that we can calculate the traing FLOPs if we know the total edge number in the dataset.
Appendix G The Model Scaling Behaviors of Node Classification and Link Prediction

### v3

@@ Appendix I How do Our Findings Connect to Compute Scaling and Chinchilla Scaling Law
where $U$ and $M$ can be assumed to contain $x$ and $y$ add-multiply operations, respectively. Suppose the total node number in the dataset is N, the total edge number is E, and the model has L layers. For node $i$, aggregating messages from its neighbors needs $x+yd_{i}$ add-multiply operations, where $d_{i}$ is the degree of node $i$. For the whole datasets, there would be $\sum_{i}x+yd_{i}=2(1+1/d_{a})yE$, where $d_{a}$ is the average degree of the dataset. For the whole model, one feed-forward pass needs $2(1+1/d_{a})yEL$ add-multiply operations. Thus, we have shown that we can calculate the training FLOPs if we know the total edge number in the dataset.
Chinchilla scaling law aims to allocate the best training set size and the model size when the compute cost is fixed [15]. The graph scaling law developed in our paper can be easily extended to calculate optimal parameter numbers or dataset size as shown below. For simplicity, we just consider Equation 4 here. But the results can be easily extended to Equation 5. For Equation 4 we have
$\epsilon=\epsilon_{\infty}+A/N^{\alpha}+B/D^{\beta}$
where N is the model size, and D is the training set size. From the discussion above and in [2], we know the FLOPs for GIN are about $6\frac{1+d_{a}}{d_{a}}EN$, where E is the number of edges and $d_{a}$ is the average degree of the training set. For fixing the compute budget $C=6\frac{1+d_{a}}{d_{a}}EN$, we can get $N_{opt}=G(\frac{d_{a}C}{6(1+d_{a})})^{a}$, $E_{opt}=G^{-1}(\frac{d_{a}C}{6(1+d_{a})})^{b}$ following [2], where $G=(\frac{\alpha A}{\beta B})^{\frac{1}{\alpha+\beta}}$, $a=\frac{\beta}{\alpha+\beta}$, $b=\frac{\alpha}{\alpha+\beta}$. We can find that there is a power-law relationship between the optimal parameters/ training set size and the FLOPs.
Appendix J The Model Scaling Behaviors of Node Classification and Link Prediction